# New bound on the Yukawa coupling from CMB

### v1

@@ Abstract
We investigate the one loop inflation stemming from the superstring theory in the braneworld scenario. The tensor to scalar ratio of the loop inflation is found inconsistent with the recent CMB results for the Yukawa coupling from the SM sector. We propose a new bound on the Yukawa coupling, , applicable to the cosmological sector from CMB. The aftermath of this new bound is explored. The present results may shed some light on the phenomenology of superstring theory and its associated phenomena.
Keywords : Inflation, Supersymmetry, String theory, CMB, Yukawa coupling.
@@
In SM particle phenomenology, the well-celebrated Yukawa coupling provides the strength of the interaction between fermions and bosons. Existing values [1, 2, 3] of the Yukawa coupling are estimated mostly from various SM particle interactions. These coupling strengths can be challenged in the inflationary models stemming from supersymmetric potentials. The superpotential and the Yukawa couplings are determined by the underlying supersymmetry (SUSY). In this work, we show that adopting the current values from the SM sector Yukawa coupling (SMYC) in the superstring theory motivated loop inflation leads to an incompatibility of the tensor to scalar ratio with recent CMB estimates. We propose a new bound on the Yukawa coupling applicable to the cosmological sector from CMB. Consequences of the newly derived Yukawa coupling limits are also discussed.
@@
The genesis of one loop inflationary model [4, 5, 6, 7, 8] is from hypersymmetric -term inflation implemented in the D3/D7 brane that reduces to a supersymmetric -term inflation by altering the flatness of the inflationary potential via one loop radiative correction. The resulting Einstein-Hilbert action on the brane is
@@
is the gauge coupling and is the Fayet-Iliopoulos term. The gauge coupling is related to the Yukawa coupling, . Here, we take the critical value of the inflaton field . We assume that inflation proceeds in the slow roll regime and express the first and second slow-roll parameters for in terms of and the e-folding number () respectively as
@@
To check the feasibility of presently available values from SMYC [3] on the loop inflation, we estimate the scalar spectral index and tensor to scalar ratio (see table 1) then compare them with recent CMB estimates (see figure 1).
From the comparative study, we can see that the tensor to scalar ratio of the loop inflation with the current bound from SMYC for the range [3] is much higher than the upper limit derived from BK18 (0.036) [14], PR 4 + BK18 (0.032) [15], LiteBIRD [16] and CMB S4 (0.004, future projected estimate) [17]. Moreover, analysis of the variation of tensor to scalar ratio with the scalar spectral index shows that all are out of the 95 and 68 CL of both Planck2018 and BK18 bounds (see figure 1). We observe that lies in 68 CL of Planck2018 but in 95 CL of BK18. Clearly, we can notice that for all , the loop inflationary scenario is disfavoured based on the current and future CMB observations with the existing values of the Yukawa coupling constant from SM sector under consideration. Since the loop inflation originates from the superstring theory, discarding the model can in turn challenge supersymmetry, string theory and quantum gravity. At the same time, it is interesting to note that lies well within the 68 CL of both Planck2018 and BK18. Therefore, at a glance, the Yukawa coupling from SM sector appears to be a viable candidate for accounting the Planck2018 and BK18 with 68 CL. However, further scrutiny shows that this value cannot be considered because it comes from a semi-classical approximation (not radiatively corrected) [3] whereas the loop inflation is a radiatively corrected model. The tensor to scalar ratio of this inflationary model can be re-estimated with the values from SMYC [1, 2, 3] other than the range . Even then the conclusion remains unaltered. This alarming situation prevents us from adopting SMYC directly into the cosmological sector. Therefore, there is a necessity to investigate the values of Yukawa coupling beyond SM interactions, particularly suitable for the braneworld cosmology. For this, the Yukawa interaction strength in the braneworld cosmological sector need to be derived. Hence, we estimate values of the Yukawa coupling constant for the one loop inflationary model from Planck2018 and BK18. The newly derived results are presented in table 2. Since these values are estimated based on CMB that comes from the cosmological observations, we call them the cosmological sector Yukawa coupling (CSYC).
@@
We study the feasibility of CSYC with the scalar spectral index and tensor to scalar ratio obtained from the joint Planck TT,TE,EE+lowE+lensing [12] and BK18+BAO [13, 14] with the marginalized 95 and 68 CL (see figure 2). The analysis shows that falls within the 68 CL of Planck2018 and 95 of BK18. But, all fall within the 95 CL of Planck2018, however, they are outside the 95 and 68 CL of BK18 bounds and 68 CL of Planck2018. Similarly, all the values of fall outside the 95 CL of both BK18 and Planck2018 hence are not feasible CSYC for the one loop inflation. Therefore, to account for the current and future estimated tensor to scalar ratio, the situation enforces us to adopt a new range of the Yukawa coupling for the one loop inflation. This suggests a new bound on the Yukawa coupling applicable in the cosmological sector from CMB which is .
@@
The one loop inflationary model originates from supersymmetric theory. Therefore, the Yukawa coupling strength associated with it can play a pivotal role in deciding its fate on the braneworld inflationary scenario. Hence, the right value of Yukawa coupling must be used so that the model is compatible with various CMB observations, especially the tensor to scalar ratio. Presently, values of the Yukawa coupling strength are estimated only from various SM sectors and exhibit a very wide range. However, selecting the coupling strength from a pool of SM sector and using it in the loop inflation leads to inconsistency of its scalar spectral index and tensor to scalar ratio compared with the respective estimates from current CMB. This raises a concern about the adoptability of Yukawa coupling from the SM sector into the cosmological sector in a straight forward manner. This circumstance demands to investigate suitable values of the Yukawa coupling constant for the viability of one loop inflation with various results of CMB. Therefore, we derived a new bound of Yukawa coupling constant from 68 CL of Planck2018 and BK18 from CMB that is applicable to cosmology. To the best of knowledge, the authors are not aware of such a bound derived on Yukawa coupling from any cosmological sector so far. The new derived constraint on the Yukawa coupling constant from CMB may play a crucial role in validating inflationary model coming from supersymmetry. Yukawa coupling from string theory, supersymmetry, quantum gravity and extra dimensions can bridge the SM and cosmological sectors.
@@
In modern times, the Yukawa coupling is estimated from the high energy particle collider experiments such as ATLAS, CMS,..etc [1, 2] and such values are used to estimate mass of quarks, leptons,..etc. Alternatively, we suggest a similar possible estimate from cosmological sector by considering the early universe phenomena through CMB. Therefore, it can provide a platform to compare the values of Yukawa coupling strength arising from the cosmological sector with the SM sector. That may lead to interesting results which in turn complements in understanding particle physics, supersymmetry, string theory, quantum gravity and cosmology. The difference in the values of the Yukawa coupling from the two distinct prominent sectors (SMYC and CSYC) arises because of the involved energy scales of interaction in the different sectors. Interestingly, SMYC estimates lie in 95 CL of ATLAS/CMS [1, 2] whereas CSYC falls in the 68 CL of Planck2018 and BK18.
Standard inflationary scale ( GeV) is much higher than currently running LHC experiments even at the TeV scale.
Moreover, earlier LHC experiments were unsuccessful in providing any evidence for SUSY. This could be due to the extremely high energy scale of SUSY than LHC. This can lead to a wrong conclusion about SUSY and related phenomena. This picture can be altered if we can get some evidence or signal from the early universe phenomena such as superstring theory driven loop inflation through a suitable cosmological probe such as CMB. We considered the one loop inflation emerging from superpotential and studied its tensor to scalar ratio and scalar spectral index with the 68 CL of both Planck2018 TT,TE,EE+lowE+lensing and BK18+BAO and demonstrated that the Yukawa coupling can exist beyond SM sector. Hence, the newly derived Yukawa coupling strength from CMB can substantiate SUSY and its related phenomena. In short, we can say that the estimated Yukawa coupling from the cosmological sector and SM sector can complement to widen or deepen the understanding so that the limitation of the LHC experiments can be overcome by looking at the cosmological phenomenology in estimating the Yukawa coupling for interactions taking place at energies higher than the LHC operating scale.
@@
The possibility of estimating the Yukawa coupling from the superstring theory based loop inflation with CMB can overcome the strong criticism against the phenomenological feasibility of string theory and its related ideas unambiguously. This may once again resurrect the wonderful idea of extra dimensions and superstring theory as a candidate for quantum gravity and can generate a renewed interest among the high energy and cosmology community. In that respect, our work takes a leap.
The study can be repeated with inflationary models arising from string theory that may support or widen or narrow the derived Yukawa coupling bounds. We hope that the current result can revive superstring and quantum gravity. The newly derived bound from CMB may be a key to open the superstring theory and associated phenomena. We are successful in deriving a new bound on the Yukawa coupling from the cosmological sector.
@@
The new bound indicates the existence of fermions with very high mass. The question is what happened to such fermions? One of the possibility is that, the high mass fermion decayed into other smaller particles. Some of them turned into DM/DE. Other possibility is that the decayed fermionβs density considerably reduced because of inflation so that we dont see them in the present universe. At the same time a completely different aspect of it is that the production of such massive fermions is unlikely to happen which implies no signal for extra dimension and superstring theory. Since the derived new bound of Yukawa coupling from CMB therefore supports the same conclusion regarding the extradimension and SUSY as that of the LHC experiment. Therefore the result of the present work may also support the LHC conclusions. There will be many consequences if this bound is not accepatable including the ruling out of loop inflation which inturn negates supersymmetry.
Acknowledgments

### v2

# Feasibility of one loop inflation in the light of CMB

@@ Abstract
The one loop inflation stemming from the superstring theory and associated Yukawa coupling arising from supersymmetric interactions is examined with CMB. The Yukawa coupling can exist beyond standard model particle physics sector. The tensor to scalar ratio of the loop inflation is found consistent with the recent CMB results for the Yukawa coupling from cosmology. The newly derived constraint on the Yukawa coupling constant may play a crucial role in validating inflationary model originating from supersymmetry and may shed some light on the formation of dark matter or dark energy. The outcomes of the study may be helpful in the phenomenological realisation of string theory.
Keywords: Supersymmetry, superstring, inflation, CMB
Anupama B11121phph19@uohyd.ac.in and P K Suresh222sureshpk@uohyd.ac.in
School of Physics, University of Hyderabad, Hyderabad, 500046, India.
1 Introduction
Inflation is considered as a period of exponential expansion in the evolution history of early universe [1]. It stretches the zero-point quantum oscillations of a homogenous scalar field called inflaton responsible for inflation. One of the important features of inflation includes the generation of scalar and tensor perturbations that respectively lead to the formation of large scale structure of the universe and the generation of primordial gravitational waves. The fractional contribution of these tensor and scalar modes can be quantified using ratio of their amplitudes, known as tensor to scalar ratio . In standard parlance, inflationary models use a single scalar field to describe the dynamics of inflation, but over the years many single and multi fields inflationary models have also been proposed [2]. The validation of these models depends primarily on the tensor to scalar ratio. The tensor to scalar ratio of most of these models is found incompatible with the estimates from various cosmological observations [3, 4, 5, 6, 7, 8, 9]. Therefore it is interesting to explore inflationary models that can possibly make the tensor to scalar ratio compatible with the current [8] and future projected estimates [10, 11].
Apart from the usual classical models, inflation can be described using supersymmetric string theory that predicts a 10-dimensional spacetime with gauge bosons and fermions [12, 13]. There are five types of superstring theories such as type-I, type-IIA, type-IIB, heterotic and heterotic . Strings can be considered as 1-dimensional analogues of higher dimensional dynamical objects called -branes. Open strings can be localised to higher dimensional membranes called D-branes by imposing Dirichlet boundary conditions to their endpoints whereas the closed strings enjoy the freedom for propagation through extra dimensions called bulk [14, 15]. Compactification of extra dimensions into the 4 dimensions of spacetime can be carried out using the models like the Kaluza-Klein model [16] and the Calabi-Yau model [17, 18]. Kaluza-Klein theory was proposed to unify electromagnetic theory and gravity in 5 dimensions before the advent of general theory of relativity. Supersymmetry (SUSY) can be accommodated in Calabi-Yau manifold without symmetry breaking and hence can be considered as the best compactification model so far. A paradigm shift in string theory came with the unification of these theories into a new supergravity M-theory with 11 dimensions [19, 20].
Efforts have been made to study inflation by incorporating supersymmetry into general theory of relativity called supergravity (SUGRA) but, are not succesful completely because the inflation can be hindered due to the production of massive gravitons [21]. Therefore the brane inflation was introduced with more than one supersymmetry (an extended supersymmetric gauge theory where is a collection of supercharges) where the matter content of the model is given by a vector multiplet and a charged hypermultiplet [12]. The ground state of brane modes is considered as inflaton and is an open string with its two ends residing on different branes. The inflation is driven by the vacuum energy density due to the initial separation of branes. Inflation occurs through the inter brane interaction and can be described by the slowly rolling inflaton field. The resulting brane motion can bring the different branes together that were initially separated [22, 23, 24]. The inflation ends as soon as the branes approach closer and reach a critical distance, i.e, when the inflaton moves away from the unstable state and occupies the supersymmetric vacuum. Several attempts have been made to compute the tensor to scalar ratio from string theory motivated inflationary models. Corresponding tensor to scalar ratio can be compared with CMB estimates and can provide a platform to test the underlying inflationary model and hence the string theory [25, 26].
In standard cosmology, the inflationary energy scale ( GeV) is much higher than the currently running large hadron collider (LHC) experiments.
Moreover, earlier LHC experiments were unsuccessful in providing any evidence for SUSY. This could be due to the extremely high energy scale of SUSY than LHC. This can lead to a wrong conclusion about SUSY and related phenomena. This picture can be altered if we can get some evidence or signal from the early universe phenomena such as superstring theory driven loop inflation through a suitable cosmological probe such as cosmic microwave background (CMB).
@@ 1 Introduction
Recently the one loop inflation gained much attention in exploring SUSY, brane inflation and Yukawa coupling. In the present work, we study the one loop corrected -term inflation which is a special case of a generic hybrid -term brane inflation in Calabi-Yau manifold in the context of type-IIB superstring theory. We show that the tensor to scalar ratio of one loop inflation is incompatible with the recent CMB estimates while using the Yukawa coupling from standard model (SM) particle physics sector. In SM particle physics phenomenology, the well celebrated Yukawa coupling provides the strength of the interaction between fermions and bosons. Existing values [27, 28, 29] of the Yukawa coupling are estimated mostly from various SM particle interactions and such values are used to estimate mass of quarks, leptons,..etc. These coupling strengths can be challenged in the inflationary models stemming from supersymmetric potentials. The superpotential and the Yukawa couplings are determined by the underlying SUSY. Hence, we derive a new bound on the Yukawa coupling arising in the supersymmetric interactions from the cosmological sector. Further, we show that the tensor to scalar ratio of loop inflation with newly derived Yukawa coupling is compatible with the recent CMB results and can be crucial in validating SUSY. Consequences of the newly derived Yukawa coupling are also discussed.
2 Loop inflation
@@
Recently the loop inflation received much attention in cosmology. The genesis of one loop inflationary model is from hypersymmetric -term inflation implemented in the D3/D7 brane that reduces to a supersymmetric -term inflation by modifying the flatness of the inflationary potential via one loop radiative correction [2, 30, 31, 32, 33, 34].
2.1 -term inflation
-term inflation is usually defined in the framework of superconformal SUSY with the auxiliary fields of the vector multiplets implemented in D3/D7 branes i.e, one end of the superstring lies on the D3 brane and the other end on D7 brane [30]. The inflation occurs in the slow roll regime where the D3 brane moves towards the D7 brane and ends with the symmetry breaking [31]. We assume the interaction of the inflaton field () with the positively and negatively charged superfields (respectively, and ) in de-Sitter spacetime with a positive cosmological constant (). The de-Sitter stage requires a soft symmetry breaking of to . This can be achieved by introducing a Fayet-Iliopoulos term (FI) (=1, 2, 3),
that can point in any direction.
The global interaction potential for a generic extended supersymmetric -term inflationary model can be written as
where is the Yukawa coupling constant and is the gauge coupling constant. The potential, Eq.(2.1), corresponds to a () supersymmetric model with the following choice of superpotential () and -term
The potential, Eq.(2.1), has two minima, non-supersymmetric unstable vacuum () through which the inflaton rolls slowly towards a supersymmetric global minimum (, Β ) achieved through gauge symmetry breaking which is essential for the end of inflation. Here, represents the critical value of the inflaton field. Inflation requires a flat direction of the potential for which one can choose and . Also, during inflation the inflaton field dominates over the other moduli fields () therefore the inflationary potential is
Inflation cannot be terminated as long as the inflationary potential remains flat. Therefore, to drive the inflationary potential towards its true vacuum, the flatness of the potential must be modified. This can be achieved by including loop corrections. Therefore the effective potential of -term inflation with loop corrections can be written as
where, , and represents the higher order corrections. A generic -term inflation is a combination of two supersymmetric inflationary models, -term inflation with chiral multiplet and -term inflation with vector multiplet respectively. But for inflation to happen smoothly, the massless -term inflation must dominate over the -term [23]. Whether -term inflation proceeds through -term or -term is decided by the value of the FI term. In the case of -term model, , and and for -term . corresponds to -term inflation in supergravity.
2.2 -term inflation
A hybrid hypersymmetric -term inflation can be reduced to a supersymmetric -term inflation by choosing under the special condition where the gauge coupling is related to the Yukawa coupling, [2, 32, 33]. -term inflation is radiatively corrected upto one loop order [32, 34]. Therefore, the global interaction potential and its one loop correction becomes
Here, the critical value of the inflaton field . The resulting Einstein-Hilbert action on the brane is
@@
By assuming the inflation proceeds in the slow roll regime, the first and second slow-roll parameters for in terms of and the e-folding number () respectively can be expressed as
@@
We study the scalar spectral index and tensor to scalar ratio of the one loop inflation with gauge coupling estimated from the presently available Yukawa coupling values from standard model particle physics [29] and is presented in Table 1. Next, we check the feasibility of this result for various parameters of the one loop inflation with the recent CMB estimates. From the comparative study (see Fig.1), it can be observed that the tensor to scalar ratio of the loop inflation with the current bound from SMYC for the range [29] is much higher than the upper limit derived from BK18 (0.036) [35], PR 4 + BK18 (0.032) [36], LiteBIRD [37] and CMB S4 (0.004, future projected estimate) [11]. Moreover, analysis of the variation of tensor to scalar ratio with the scalar spectral index shows that all are out of the 95 and 68 CL of both Planck2018 and BK18 bounds (see Fig.1). We can see that lies in 68 CL of Planck2018 but in 95 CL of BK18. Clearly, it can be noticed that for all , the loop inflationary scenario is disfavoured based on the current and future projected CMB observations with the existing values of the Yukawa coupling constant from SM particle physics sector under consideration.
At the same time, it is interesting to note that lies well within the 68 CL of both Planck2018 and BK18. Therefore, at a glance, the Yukawa coupling from SM particle physics sector appears to be a viable candidate for accounting the Planck2018 and BK18 with 68 CL. However, further scrutiny shows that this value cannot be considered because it comes from a semi-classical approximation (not radiatively corrected) [29] whereas the loop inflation is a radiatively corrected one. The tensor to scalar ratio of this inflationary model can be re-estimated with the values from SMYC [27, 28, 29] other than the range . Even then the conclusion remains unaltered. This alarming situation prevents us from adopting SMYC directly into the cosmological sector. Since the loop inflation originates from the superstring theory, discarding the model can in turn challenge supersymmetry, string theory and quantum gravity.
@@
3 Yukawa coupling from CMB
There is a necessity to investigate the values of Yukawa coupling beyond SM particle physics interactions, particularly suitable for the string motivated inflationary model. For this, the Yukawa interaction strength in the string driven cosmology need to be derived. Hence, using Eq.(13) and Eq.(14) we estimate values of the Yukawa coupling constant for the one loop inflationary model from Planck2018 and BK18. The newly derived results are presented in Table 2. Since these values are estimated based on CMB that comes from the cosmological observations, we call them the cosmological sector Yukawa coupling (CSYC).
We study the feasibility of CSYC with the scalar spectral index and tensor to scalar ratio obtained from the joint Planck TT,TE,EE+lowE+lensing [8] and BK18+BAO [35, 38] with the marginalized 95 and 68 CL (see Fig.2). The analysis shows that falls within the 68 CL of Planck2018 and 95 of BK18. But, all fall within the 95 CL of Planck2018, however, they are outside the 95 and 68 CL of BK18 bounds and 68 CL of Planck2018. Similarly, all the values of fall outside the 95 CL of both BK18 and Planck2018 hence are not feasible CSYC for the one loop inflation. Therefore, to account for the current and future estimated tensor to scalar ratio, the situation enforces us to adopt a new range of the Yukawa coupling for the one loop inflation. This suggests a new bound on the Yukawa coupling applicable in the cosmological sector from CMB which is .
@@ 3 Yukawa coupling from CMB
The one loop inflationary model originates from supersymmetric theory. Therefore, the Yukawa coupling strength associated with it can play a pivotal role in deciding its fate on the string theory based inflationary scenario. Hence, the right value of Yukawa coupling must be used so that the model is compatible with various CMB observations, especially the tensor to scalar ratio. As we have mentioned, presently, values of the Yukawa coupling strength are estimated only from various SM particle physics sectors and exhibit a very wide range. However, selecting the coupling strength from a pool of SM particle physics sector and using it in the loop inflation leads to inconsistency of its scalar spectral index and tensor to scalar ratio compared with the respective estimates from current CMB. This raises a concern about the adoptability of Yukawa coupling from the SM particle physics sector into the cosmological sector in a straightforward manner. This circumstance demands to investigate suitable values of the Yukawa coupling constant for the viability of one loop inflation with various results of CMB. Therefore, we derived a new bound on the Yukawa coupling constant from 68 CL of Planck2018 and BK18 from CMB that is found applicable to cosmology. To the best of knowledge, the authors are not aware of such a bound derived on Yukawa coupling from any cosmological sector so far. The newly derived constraint on the Yukawa coupling constant from CMB may play a crucial role in validating inflationary model coming from supersymmetry. Yukawa coupling from string theory, supersymmetry, quantum gravity and extra dimensions can bridge the SM particle physics and cosmological sectors.
4 Discussion and conclusion
It is intriguing to note that SUSY is not established observationally till date, even at various high energy accelerator experiments such as LHC,..etc. Therefore, an alternate suggestion is to explore SUSY with inflationary model originating from supersymmetry in the light of recent CMB observations.
Upon examining the superstring theory based loop inflation with the recent CMB observations, it is found that the model is disfavoured due to the adoption of SMYC in the associated fermionic and bosonic interactions. In SMYC, the Yukawa coupling is estimated from the high energy particle collider experiments such as ATLAS, CMS,..etc [27, 28]. Another possibility is estimating the Yukawa coupling from cosmology by considering the one loop inflation through CMB. We investigated the one loop inflation emerging from superpotential and studied its tensor to scalar ratio and scalar spectral index with the 68 CL of both Planck2018 TT,TE,EE+lowE+lensing and BK18+BAO and demonstrated that the Yukawa coupling is possible beyond SM particle physics sector. The newly derived bound from CMB
can substantiate SUSY and its related phenomena. In short, we can say that the estimated Yukawa coupling from the cosmological sector and SM particle physics sector can complement to widen or deepen the understanding so that the limitation of the LHC experiments can be overcome by looking at the cosmological phenomenology in estimating the Yukawa coupling for interactions taking place at energies higher than the LHC operating scale.
@@ 4 Discussion and conclusion
That may lead to interesting results which in turn complement in understanding particle physics, supersymmetry, string theory, quantum gravity and cosmology. Therefore, it can provide a platform to compare the values of Yukawa coupling strength arising from the cosmological sector with the SM particle physics sector. The difference in the values of the Yukawa coupling from the two distinct prominent sectors (SMYC and CSYC) arises because of the involved energy scales of interaction in the different sectors. Interestingly, SMYC estimates lie in 95 CL of ATLAS/CMS [27, 28] whereas CSYC falls in the 68 CL of Planck2018 and BK18.
@@
The outcomes of the present work may be useful in overcoming the strong criticism against the phenomenological realisation of string theory and its related ideas unambiguously. This may once again resurrect the wonderful idea of extra dimensions and superstring theory as a candidate for quantum gravity and can generate a renewed interest among the high energy and cosmology community. In that respect, our work takes a leap.
@@
The new bound indicates the existence of fermions with very high mass. The question is what could have happened to such fermions? One of the possibilities is that the high mass fermion decayed into other smaller particles. Some of them turned into dark matter or dark energy. Other possibility is that the decayed fermionβs density considerably reduced because of inflation so that we donβt see them in the present universe. At the same time a completely different aspect of it is that the production of such massive fermions is unlikely to happen which implies no signal for extra dimension and superstring theory. Since the derived new bound of Yukawa coupling from CMB therefore supports the same conclusion regarding the extradimension and supersymmetry as that of the LHC experiment. Therefore the result of the present work may also support the LHC conclusions. There will be many consequences if this bound is not acceptable including the ruling out of loop inflation which in turn negates supersymmetry. The study can be repeated with inflationary models arising from string theory that may support or widen or narrow the derived Yukawa coupling bounds. We hope that the current result can revive superstring and quantum gravity. Further, the new bound on the Yukawa coupling may be re-examined with the upcoming CMB missions.
Acknowledgement